OpenStudy (anonymous):

A particle experiences a constant force after starting from rest. If it travels a distance x in the first two seconds and a distance y in the next two seconds, then (Marks:1) y = 4x y = 3x y = 2x y = x

OpenStudy (mos1635):

 x=1/2 a 2^2=>x=2a z=1/2 a 4^2 =>z=8a y=z-x=6a